# Zero shot Generalization of Vision-Based RL Without Data Augmentation

### v1

@@ 4.2 Association
Figure 3 shows a comparison of using SAC with QLAE vs with BioAE Whittington et al. (2023), another disentanglement method that uses biologically inspired constraints and achieves comparable results on disentanglement benchmarks (Hsu et al., 2023). BioAE achieves strong initial performance on the two evaluation environments, but slowly degrades over the course of training. We suspect both models overfit to the training environment, but QLAE is capable of zero-shot mapping OOD latent variables to in-distribution values. To see the similarity, we first present the general framework described in Millidge et al. (2022) of a universal Hopfield network that all feedforward associative memory networks in the literature can be factorized as:
@@
$P$ is a projection, sep is a separation function, and sim is a similarity function between the stored memories $X$ and query $\xi$. While $P$ is originally described as a projection matrix, we extend the definition of $P$ to be any function that projects $X$ and $\xi$ into a shared embedding space. In this case, equation 4 can be interpreted as follows: $f_{\theta}$ is the projection function that projects high-dimensional images o into an embedding space shared by the scalar codebooks $Z$, which can be interpreted as predetermined memories. The closest memory is recovered by computing the $L_{1}$ distance, which serves as the similarity function, and $sep$ is the argmin function. Through this view, we can rewrite the latent dynamics of QLAE in many ways, perhaps exchanging the similarity function for $L_{2}$ distance or dot product, changing the separation function, etc., as long as it follows the framework of 6. Since $Z$ is a product of scalar codebooks, $L_{1}$ distance remains an appropriate choice for the similarity function. Instead, we augment the latent dynamics with a Softmax separation function as follows:
@@
[…]
In the limit, as $\beta$ goes to infinity, we achieve maximum separation between memories and recover equation 4. In practice, we choose a large value for $\beta$ such that we retrieve one scalar from each codebook, as originally intended, although our method works well with smaller values of $\beta$ (see appendix Section A.4 for additional results). Since large $\beta$ values can cause downstream gradients to vanish, we find that keeping the commitment loss from equation 5 helps keep the outputs of the encoder close to the values of the latent model. However, we do not optimize the codebook towards the encoder outputs i.e., we omit $\mathcal{L}_{\text{quantize}}$. This can be interpreted as having a set of task-optimized memories that the encoder must learn to map to under the Hopfield interpretation. Our final objective for ALDA is as follows:
$J(ALDA)$ $=\mathcal{L}_{\text{commit}}+\mathcal{L}_{\text{reconstruct}}$ (8)
$=\mathbb{E}_{\textbf{o}_{t}\sim\mathcal{D}}\Big{[}\Big{.}||\text{StopGradient}(f_{\theta}(\textbf{o}))-\left[\text{Softmax}(-\beta L_{1}(f_{\theta}(o),V))\odot V\right]||_{2}^{2}$
@@
where the last two terms are weight-decay on the encoder and decoder parameters controlled by $\lambda_{\theta}\text{ and }\lambda_{\phi}$, respectively. Observations collected by the policy are stored in a replay buffer $\mathcal{D}$, from which batches are randomly sampled to train ALDA. We observe that this formulation considerably improves training and evaluation performance on the "color hard" environment and to a degree, on the DistractingCS environment, as shown in Figure 4.
The obvious question is, how do we set the dimensionality of $z_{d}$, which should 1:1 correspond to the number of sources $n_{s}$ if the number of sources is unknown? While there is no rigorous method to derive $|z_{d}|$ at this time, we empirically found that setting $|z_{d}|$ to within the ballpark of the size of the observation spaces for the proprioceptive, state-based versions of the tasks seemed to work well. $|z_{d}|$ is set to 12 for all reported tasks.
@@ 5 Experiments
We compare against several baselines that together represent the full range of different learning paradigms in the literature that attempt to elicit zero-shot generalization. DARLA (Higgins et al., 2017b) is, to the best of our knowledge, the only other algorithm that attempts to learn a disentangled representation of the image distribution towards zero-shot generalization of vision-based RL. SAC+AE (Yarats et al., 2021b) uses a deterministic autoencoder with an auxiliary reconstruction objective and strong regularization that demonstrates decent zero-shot generalization capability. RePo (Zhu et al., 2023) is a model-based RL algorithm that learns a task-centric latent representation immune to background distractors. Finally, SVEA (Hansen et al., 2021) is an off-policy RL algorithm that improves training stability and performance of off-policy RL under data augmentation. As in their paper, we use the random overlay augmentation for SVEA, where images sampled from the Places (Zhou et al., 2017) dataset of 10 million images are overlayed during training. The training curves and evaluation on "color hard" and DistractingCS are presented in Figure 5.
Excluding SVEA, ALDA outperforms all baselines on both distribution shift environments. ALDA also maintains stability and high performance on the training environment, despite the disentanglement auxiliary objective and extremely strong weight decay ($\lambda_{\theta},\lambda_{\phi}=0.1$) on the encoder and decoder. We do not expect to outperform SVEA since it uses additional data sampled from a dataset of 1.8 million diverse real-world scenes, likely putting the training data within the support of the data distributions of the evaluation environments. Nevertheless, ALDA performs comparably and, in some cases, is equal to SVEA despite only seeing images from the original task. Performance degrades severely for all algorithms on the DistractingCS environment. We suspect that, in addition to the already difficult task of ignoring the background video, camera shaking affects the implicitly learned dynamics, and thus, additional finetuning may be unavoidable for this task. Still, ALDA performs better than all baselines excluding SVEA on Distracting CS as well, even matching the performance of SVEA on cartpole balance and finger spin.
@@
Figure 5: Top row: performance on the training environment. Middle and bottom rows are evaluation results on the "color hard" and DistractingCS evaluation environments, respectively. Average of 5 seeds, shaded region represents a 95% confidence interval.
@@
Figure 6: Visualization of different latent perturbations. Top: Traversal of select latents for the standard training environment. Bottom: Traversal of select latents when training directly on the color hard environment. Latents that encode for distractor variables (e.g., color) seemingly do not simultaneously encode for task-relevant variables (e.g., agent dynamics). Visualizations of all latent traversals can be found in Appendix section A.3.
@@
The disentangled representation learning field primarily uses toy datasets where the ground truth sources of the data distribution are known. Therefore, all disentanglement metrics we are aware of require knowing the sources, making it difficult to quantitatively evaluate disentanglement performance on DMControl. In the absence of any quantitative disentanglement metrics, we opt to show empirical evidence of disentanglement in our model, presented in Figure 6. In this experiment, we encode an observation o into the disentangled latent representation $z_{d}=l_{\psi}(f_{\theta}(\textbf{o}))$. We pick a latent variable $z_{d_{i}}$, traverse it while holding all other latent variables fixed, and decode the resulting latent vectors into reconstructed observations $\textbf{o}^{\prime}=g_{\phi}(z_{d}^{\prime})$. We find that each latent tends to learn information about a single aspect of the robot, for example, the orientation of the torso or the rotation of the left/right leg. We also find empirical evidence that ALDA does not discard task-irrelevant information, but rather encodes it separately from task-relevant latent variables when training ALDA directly on the color hard environment.
@@ 6 Discussion
As stated previously, the disentanglement problem by extension of nonlinear ICA is underdetermined, so there are many ways the latent space may factorize, perhaps by representing the sky and background with one latent or by separating them into two latents, etc. Given that both the task and reconstruction gradients of the critic/decoder affect the latent model/encoder, an interesting scientific and philosophical implication is that the model is potentially biased towards a disentangled representation that is useful, although there is no way to quantitatively or qualitatively show such a result at this time. Nevertheless, it remains an interesting line of further investigation from a scientific standpoint, and perhaps, philosophically, says something about whether the question "What is the ground truth factorized representation that best explains the data?" is even the right question to ask.
RL agents deployed in the real world must constantly adapt to changing environmental conditions. Much of the variance can be captured with a sufficiently large dataset. However, there remains a portion of the distribution containing every possible edge case and unaccounted-for variation, commonly referred to as "the long tail," that remains elusive because it is prohibitively expensive to account for every possible variation. Unfortunately, these uncaptured variations are frequent enough due to the ever-changing dynamical nature and complexity of the real world that deploying agents in the real world remains challenging. Therefore, it seems the case that data augmentation techniques, collecting massive datasets, and the like are not sufficient to develop generalist agents capable of adaptation the way humans and other animals are. That’s not to say that data isn’t important or a fundamental ingredient to training machine learning models. In fact, the method proposed in this paper can also utilize data augmentation techniques presented in prior work, as it should. Instead, our proposition is that if a data-driven model can generalize better with less data, then it will scale better with more data.
@@
7 Acknowledgements
@@ A.1 Proof
Preliminaries. We reintroduce the nonlinear ICA problem formulation here for the reader’s reference. There are $n_{s}$ nonlinear independent variables $\textbf{s}=s_{1},...,s_{n_{s}}$ that are considered the "ground truth" sources of variation of the data distribution. We assume there exists a data-generating model that maps sources to images:
@@
$cov(\hat{s_{i}},\hat{s_{j}}|z_{k})=0\text{ }\forall s_{i}\in D,s_{j}\in E,i\in[1,k],j\in[k+1,n_{s}].$ (12)
To see why this must be the case, suppose that the covariance is nonzero and suppose that we perturb $s_{j}\in E$ to $s_{j}^{\prime}$, giving us a new observation $\textbf{o}^{\prime}=g(\textbf{s}^{\prime})$. Since $s_{j}$ is task-irrelevant, $\textbf{z}^{\prime}=f_{\theta}(\textbf{o}^{\prime})=f_{\theta}(h(\textbf{o},v))$ for some $v\in\mathcal{T}$. If $\textbf{z}^{\prime}$ is a function of $h$, then by equations 9 and 11, $Q^{*}(\textbf{z}^{\prime},a)=Q^{*}(\textbf{z},a)$. However, if $cov(\hat{s_{i}},\hat{s_{j}}|z_{k})\neq 0$ for some $_{k}\in\textbf{z}$ z, then $\hat{s_{i}}^{\prime}\in\textbf{z}^{\prime}\neq\hat{s_{i}}\in\textbf{z}$. Since $\hat{s_{i}}\in D$, $Q^{*}(\textbf{z}^{\prime},a)\neq Q^{*}(\textbf{z},a)$, which is a contradiction. Therefore, the conditional covariance between any $\hat{s_{i}}\in D$ and any $\hat{s_{j}}\in E$ for any given $z_{k}$ must be zero, which implies that the approximations of task-relevant and task-irrelevant sources in z are disentangled.
@@ A.2 Latent Trajectory Visualizations
Given the low dimensionality of the disentangled latent representations and the fact that they are disentangled, we hypothesize that their trajectories through time may correspond to trajectories of proprioceptive state variables such as rigid-body positions and orientations. We visualize the trajectories of individual latents through time for a single episode alongside the trajectories of several proprioceptive state variables in Figure 7. Unfortunately, the mappings of sources to the disentangled latent space likely do not correspond 1:1 with the proprioceptive state, given that the learned mappings are arbitrary and not unique. However, upon visual inspection, we find that latent trajectories through time exhibit oscillatory behavior patterns similar to that of rigid body orientations from the state representation. Recovering all or parts of the proprioceptive state representation via unsupervised learning from high-dimensional data is an interesting future research direction.
@@ A.3 Additional Latent Traversal Plots
Latent traversals for the other reported DMControl tasks are presented here. We also visualize the latent traversals of ALDA trained directly on the color hard environment. We find that the latent traversals for cartpole balance are more discontinuous than on other tasks. One reason for this might be the lack of balanced data and data diversity of the cartpole replay buffer. The (near) optimal policy is achieved quite early on, after which most images collected are of the cartpole upright and roughly in the same $x$-position. The lack of data diversity likely makes it difficult to learn a representation in which the latent traversals are more continuous / physically plausible. Interestingly, this phenomenon does not seem to affect performance on the "color hard" evaluation environment, although we suspect there are performance gains to be had on DistractingCS if the latent interpolations were smoother. We leave an investigation into the effects of data balancing and data diversity on downstream generalization performance as future work.
Figure 8: Latent traversals of the disentangled latent vector when training ALDA directly on the "color hard" environment.

### v2

@@ 4.2 Association
Figure 3 shows a comparison of using SAC with QLAE vs with BioAE (Whittington et al., 2023), another disentanglement method that uses biologically inspired constraints and achieves comparable results on disentanglement benchmarks (Hsu et al., 2023). This comparison allows us to isolate and analyze how the associative latent dynamics contributes to generalization, since BioAE does not incorporate any associative mechanism while QLAE implicitly does. BioAE achieves strong initial performance on the two evaluation environments, but slowly degrades over the course of training. We suspect both models overfit to the training environment, but QLAE is capable of zero-shot mapping OOD latent variables to in-distribution values. To understand how QLAE implicitly incorporates an associative mechanism, we first present the general framework described in Millidge et al. (2022) of a universal Hopfield network that all feedforward associative memory networks in the literature can be factorized as:
@@
$P$ is a projection, sep is a separation function, and sim is a similarity function between the stored memories $X$ and query $\xi$. In this case, equation 3 can be interpreted as follows: we compute similarity scores via $L_{1}$ distance between $f_{\theta}(x)$ and $\text{StopGradient}(z_{d})$, and the separation function is the argmin operator. Since argmin retreives the desired value from the codebook, the projection function $P$ in this case is the identity function. Through this view, we can rewrite the latent dynamics of QLAE in many ways, perhaps exchanging the similarity function for $L_{2}$ distance or dot product, changing the separation function, etc., as long as it follows the framework of 5. Since $Z$ is a product of scalar codebooks, $L_{1}$ distance remains an appropriate choice for the similarity function. Instead, we augment the latent dynamics with a Softmax separation function as follows:
@@
where $\beta$ is a scalar temperature parameter. Equation 6 can be interpreted in two ways. From an associative memory perspective, attention-based Hopfield models apply Softmax to separate the local minima (stored memories) on the energy landscape, where $\beta$ controls the degree of separation, and so we’ve recovered the modern Hopfield memory retrieval dynamics. From a purely mathematical perspective, we have what resembles the Gumbel-Softmax categorical reparameterization Jang et al. (2017), although we do not perform any sampling in our method. This lends a novel view on attention-based Hopfield networks – models with a high-temperature parameter can be interpreted as classifiers over $|\textbf{X}|$ classes, where $|\textbf{X}|$ is the number of stored memories whose local minima are well separated on the energy landscape.
In the limit, as $\beta$ goes to infinity, we achieve maximum separation between memories and recover equation 3. In practice, we choose a large value for $\beta$ such that we retrieve one scalar from each codebook, as originally intended, although our method works well with smaller values of $\beta$ (see appendix Section A.4 for additional results). Since large $\beta$ values can cause downstream gradients to vanish, we find that keeping the commitment loss from equation 4 helps keep the outputs of the encoder close to the values of the latent model. However, we do not optimize the codebook towards the encoder outputs i.e., we omit $\mathcal{L}_{\text{quantize}}$. This can be interpreted as having a set of task-optimized memories that the encoder must learn to map to under the Hopfield interpretation. Our final objective for ALDA is then simply $\mathcal{L}_{\text{commit}}+\mathcal{L}_{\text{reconstruct}}$:
$J(ALDA)$ $=\mathbb{E}_{\textbf{o}_{t}\sim\mathcal{D}}\Big{[}\Big{.}||f_{\theta}(\textbf{o})-\text{StopGradient}($ (7)
$\left[\text{Softmax}(-\beta L_{1}(f_{\theta}(o),V))\odot V\right]||_{2}^{2})$
@@
where the last two terms are weight-decay on the encoder and decoder parameters controlled by $\lambda_{\theta}\text{ and }\lambda_{\phi}$, respectively. Observations collected by the policy are stored in a replay buffer $\mathcal{D}$, from which batches are randomly sampled to train ALDA. We observe that this formulation considerably improves training and evaluation performance on the color hard environment and to a degree, on the DistractingCS environment, as shown in Figure 4. The obvious question is, how do we set the dimensionality of $z_{d}$, which should 1:1 correspond to the number of sources $n_{s}$ if the number of sources is unknown? While there is no rigorous method to derive $|z_{d}|$ at this time, we empirically found that setting $|z_{d}|$ to within the ballpark of the size of the observation spaces for the proprioceptive, state-based versions of the tasks seemed to work well. $|z_{d}|$ is set to 12 for all reported tasks.
@@ 5 Experiments
Figure 5: Top row: performance on the training environment. Middle and bottom rows are evaluation results on the ”color hard” and DistractingCS evaluation environments, respectively. Average of 5 seeds, shaded region represents a 95% confidence interval.
@@
Figure 6: Latent traversals. Each row corresponds to holding all but one latent variable fixed, interpolating it from min to max value in its codebook, and decoding the resulting vector into an image. Top: Traversal of select latents for the standard training environment. Bottom: Traversal of select latents when training directly on the color hard environment. Latents that encode for distractor variables (e.g., color) seemingly do not simultaneously encode for task-relevant variables (e.g., agent dynamics). Visualizations of all latent traversals can be found in Appendix section A.3.
We compare against several baselines that together represent the full range of different learning paradigms in the literature that attempt to elicit zero-shot generalization. DARLA (Higgins et al., 2017b) is, to the best of our knowledge, the only other algorithm that attempts to learn a disentangled representation of the image distribution towards zero-shot generalization of vision-based RL. SAC+AE (Yarats et al., 2021b) uses a deterministic autoencoder with an auxiliary reconstruction objective and strong regularization that demonstrates decent zero-shot generalization capability. RePo (Zhu et al., 2023) is a model-based RL algorithm that learns a task-centric latent representation immune to background distractors. Finally, SVEA (Hansen et al., 2021) is an off-policy RL algorithm that improves training stability and performance of off-policy RL under data augmentation. As in their paper, we use the random overlay augmentation for SVEA, where images sampled from the Places (Zhou et al., 2017) dataset of 10 million images are overlayed during training. The training curves and evaluation on ”color hard” and DistractingCS are presented in Figure 5.
Excluding SVEA, ALDA outperforms all baselines on both distribution shift environments. ALDA also maintains stability and high performance on the training environment, despite the disentanglement auxiliary objective and extremely strong weight decay ($\lambda_{\theta},\lambda_{\phi}=0.1$) on the encoder and decoder. We do not expect to outperform SVEA since it uses additional data sampled from a dataset of 1.8 million diverse real-world scenes, likely putting the training data within the support of the data distributions of the evaluation environments. Nevertheless, ALDA performs comparably and, in some cases, is equal to SVEA despite only seeing images from the original task. Furthermore, we find that ALDA can even outperform SVEA if SVEA is trained with different data augmentations not including image overlays from the Places dataset, shown in Appendix Section A.11. Performance degrades severely for all algorithms on the DistractingCS environment. We suspect that, in addition to the already difficult task of ignoring the background video, camera shaking affects the implicitly learned dynamics, and thus, additional finetuning may be unavoidable for this task. Still, ALDA performs better than all baselines excluding SVEA on Distracting CS as well, even matching the performance of SVEA on cartpole balance and finger spin.
@@
The disentangled representation learning field primarily uses toy datasets where the ground truth sources of the data distribution are known. Therefore, all disentanglement metrics we are aware of require knowing the sources, making it difficult to quantitatively evaluate disentanglement performance on DMControl. In the absence of any quantitative disentanglement metrics, we opt to show empirical evidence of disentanglement in our model, presented in Figure 6. In this experiment, we encode a random batch of observations o into the disentangled latent representation $z_{d}=l_{\psi}(f_{\theta}(\textbf{o}))$. We pick one latent dimension $z_{d_{i}}$ of the latent vector and interpolate it from its minimum to maximum value in its codebook while holding all other latent variables fixed. The resulting perturbed latents are decoded using ALDA’s learned decoder $\textbf{o}^{\prime}=g_{\phi}(z_{d}^{\prime})$ and presented as a row of images. We find that each latent tends to learn information about a single aspect of the robot, for example, the orientation of the torso (middle row of the top image) or left hip joint angle (top row of the top image). From the bottom image in Figure 6, we can see that ALDA encodes scene color information separately from robot dynamics when trained directly on the color hard environment. Unlike methods that learn invariant representations, ALDA does not discard task-irrelevant information, but rather encodes all variables separately, resulting in an implicit separation between task-relevant and task-irrelevant variables. Further empirical analysis on disentanglement of the image observations and what the latent dimensions learn to represent are presented in A.2.
@@ 6 Discussion
As stated previously, the disentanglement problem by extension of nonlinear ICA is underdetermined, so there are many ways the latent space may factorize, perhaps by representing the sky and background with one latent or by separating them into two latents, etc. Given that both the task and reconstruction gradients of the critic/decoder affect the latent model/encoder, an interesting scientific and philosophical implication is that the model is potentially biased towards a disentangled representation that is useful, although there is no way to quantitatively or qualitatively show such a result at this time. Nevertheless, it remains an interesting line of further investigation from a scientific standpoint, and perhaps, philosophically, says something about whether the question ”What is the ground truth factorized representation that best explains the data?” is even the right question to ask.
RL agents deployed in the real world must constantly adapt to changing environmental conditions. Much of the variance can be captured with a sufficiently large dataset. However, there remains a portion of the distribution containing every possible edge case and unaccounted-for variation, commonly referred to as ”the long tail,” that remains elusive because it is prohibitively expensive to account for every possible variation. Unfortunately, these uncaptured variations are frequent enough due to the ever-changing dynamical nature and complexity of the real world that deploying agents in the real world remains challenging. Therefore, it seems the case that data augmentation techniques, collecting massive datasets, and the like are not sufficient to develop generalist agents capable of adaptation the way humans and other animals are. That’s not to say that data is not important or a fundamental ingredient to training machine learning models. In fact, the method proposed in this paper scales with more data as with prior works that leverage data augmentation techniques. Instead, our proposition is that if a data-driven model can generalize better with less data, then it will scale better with more data.
@@
Impact Statement
Agents are rapidly being integrated into our society across many different dimensions simultaneously, from web-based LLM agents to driverless cars, to household robotics. A general trend is that as models become larger and are trained on vast datasets, the interpretability becomes diminished, which is problematic for AI safety research, especially in critical tasks as autonomous driving. It is also generally the case that methods which are more interpretable by design are interpretable at the expense of performance when compared to fully end-to-end systems without safety constraints. We believe that encouraging learning disentangled representations in our models not only enables greater interpretability and facilitates AI safety research, but does so while improving generalization rather than hindering it, and is thus a promising avenue of further research.
Acknowledgements
@@ A.1 Proof
Preliminaries. We reintroduce the nonlinear ICA problem formulation here for the reader’s reference. There are $n_{s}$ nonlinear independent variables $\textbf{s}=s_{1},...,s_{n_{s}}$ that are considered the ”ground truth” sources of variation of the data distribution. We assume there exists a data-generating model that maps sources to images:
@@
[…]
To see why this must be the case, suppose that the covariance is nonzero and suppose that we perturb $s_{j}\in E$ to $s_{j}^{\prime}$, giving us a new observation $\textbf{o}^{\prime}=g(\textbf{s}^{\prime})$. Since $s_{j}$ is task-irrelevant, $\textbf{z}^{\prime}=f_{\theta}(\textbf{o}^{\prime})=f_{\theta}(h(\textbf{o},v))$ for some $v\in\mathcal{T}$. If $\textbf{z}^{\prime}$ is a function of $h$, then by equations 8 and 10, $Q^{*}(\textbf{z}^{\prime},a)=Q^{*}(\textbf{z},a)$. However, if $cov(\hat{s_{i}},\hat{s_{j}}|z_{k})\neq 0$ for some $z_{k}\in\textbf{z}$, then $\hat{s_{i}}^{\prime}\in\textbf{z}^{\prime}\neq\hat{s_{i}}\in\textbf{z}$. Since $\hat{s_{i}}\in D$, $Q^{*}(\textbf{z}^{\prime},a)\neq Q^{*}(\textbf{z},a)$, which is a contradiction. Therefore, the conditional covariance between any $\hat{s_{i}}\in D$ and any $\hat{s_{j}}\in E$ for any given $z_{k}$ must be zero, which implies that the approximations of task-relevant and task-irrelevant sources in z are disentangled.
@@ A.2 Latent Trajectory Visualizations
Given the low dimensionality of the disentangled latent representations and the fact that they are disentangled, we hypothesize that their trajectories through time may correspond to trajectories of proprioceptive state variables such as rigid-body positions and orientations. We visualize the trajectories of individual latents through time for a single episode alongside the trajectories of several proprioceptive state variables in Figure 7. Unfortunately, the mappings of sources to the disentangled latent space likely do not correspond 1:1 with the proprioceptive state, given that the learned mappings are arbitrary and not unique. However, upon visual inspection, we find that latent trajectories through time exhibit oscillatory behavior patterns similar to that of rigid body orientations from the state representation. Given that the representation is task optimized and that the agent is the only dynamic object in the scene, we suspect that the model learns to disentangle and primarily represent the agent’s rigid bodies and their orientations. Recovering all or parts of the proprioceptive state representation via unsupervised learning from high-dimensional data is an interesting future research direction.
@@ A.3 Additional Latent Traversal Plots
Latent traversals for the other reported DMControl tasks are presented here. We also visualize the latent traversals of ALDA trained directly on the color hard environment. We find that the latent traversals for cartpole balance are more discontinuous than on other tasks. One reason for this might be the lack of balanced data and data diversity of the cartpole replay buffer. The (near) optimal policy is achieved quite early on, after which most images collected are of the cartpole upright and roughly in the same $x$-position. The lack of data diversity likely makes it difficult to learn a representation in which the latent traversals are more continuous / physically plausible. Interestingly, this phenomenon does not seem to affect performance on the ”color hard” evaluation environment, although we suspect there are performance gains to be had on DistractingCS if the latent interpolations were smoother. We leave an investigation into the effects of data balancing and data diversity on downstream generalization performance as future work.
Figure 8: Latent traversals of the disentangled latent vector when training ALDA directly on the ”color hard” environment.